OpenStudy (anonymous):

Can someone turn the equation f(x)=-1.91x^2-9.55x-7.64 from standard form into general form of a quadratic equation?

OpenStudy (anonymous):

Looks like a quadratic to me. The highest power of x is 2. And it can be factored. -1.91x^2-9.55x-7.64 = -1.91 (1.+1. x) (4.+1. x)

OpenStudy (anonymous):

is that general form?

OpenStudy (anonymous):

@iambatman @IMStuck @Nurali @Hero @ganeshie8 @Conqueror @campbell_st

OpenStudy (campbell_st):

the general form of a quadratic equation is \[y = ax^2 + bx + c...or.... f(x) = ax^2 + bx + c\] your equation is in general form a = -1.91 b = -9.55 and c = 7.64

OpenStudy (anonymous):

then what is standard form

OpenStudy (campbell_st):

you're equation already is in standard form....

OpenStudy (campbell_st):

there is a standard vertex form....

OpenStudy (anonymous):

what would that be @campbell_st

OpenStudy (anonymous):

can someone turn the equation into \[y=a(x-h)^2+k\]

OpenStudy (anonymous):

@iGreen @Compassionate @ganeshie8 @IMStuck @thomaster @phi